OpenStudy (anonymous):

a piggy bank contained $48 which consisted of $2, $1 and 50c coins. The number of $1 was twice the number of $2 coins and the number of 50c coins was twice the number of $1 coins. How many of each coin were there?

OpenStudy (anonymous):

ones=2twos cents=2ones 48=2twos+ones+.5cents 48=2twos+2twos+4twos 48=8twos twos=6 ones=2*6 ones = 12 cents=2*12 cents=24